You throw a ball with a mass of 0.40 kg against a brick wall. It hits the wall moving horizontally to the left at -30m/s and rebounds horizontally to the right at 20m/s. Find the impulse of the net force on the ball during its collision with the wall.


1
Expert's answer
2021-12-06T16:42:05-0500
"\Delta p=m(v_f-v_i),""\Delta p=0.4\ kg\times(20\ \dfrac{m}{s}-(-30\ \dfrac{m}{s}))=20\ \dfrac{kg\times m}{s}."